A restaurant manager classifies customers as regular, occasional, or new, and finds that of
all customers 25%, 56% and 19% respectively, fall into these categories. The manager
found that nshima was ordered by 70% of the regular customers, by 50% of the occasional,
and 30% of the new customers.
i. What is the probability that a randomly chosen customer orders nshima?
ii. If nshima is ordered, what is the probability that the person ordering is a regular
customer?
iii. If nshima is ordered, what is the probability that the person ordering is an occasional
customer?
iv. If nshima is ordered, what is the probability that the person ordering is a new
customer?

Answer by robertb(5830)



i. 0.25*0.70 + 0.56*0.50 + 0.19*0.30 = 7/40 + 7/25 + 57/1000 = 64/125

ii. (7/40)/(64/125) = 175/512 ~0.342, to 2 d.p.

From part ii, you should be able to see the approach for the solution of parts iii and iv.